Is 3 square root 34 rational or irrational?
Answer by RAY100(1637)

3sqrt 34 = 17.493,,,,,which is irrational
,
remember rational can be expressed as a/b
,
sqrt that are not perfect squares are a major group if irrational numbers
,
for decimals remember, termial decimals are rational, ie, .25 =1/4,,,.375 = 3/8,, etc
,
infinite , but repeating decimals are rational, .3333rptg =1/3 etc
,
infinite but not repeating, are irrational,,,,(cannot fit a/b form) this is what we have above
,
keep up good work